OpenStudy (anonymous):

Use Cauchy-Hadamard to find the Radius of convergence and the center of the series. ((-1)^n)/((2n)!) * (z- 1/4 pi)^2n

OpenStudy (anonymous):

Thing is , my answer is absolutely different from the one I am supposed to get. I get infinity as radius sure enough, but my radius is pi/4 when it should be pi/2

OpenStudy (anonymous):

So, if I understand this theorem correctly you need to find the actual c_n? To evaluate the lim sup?

OpenStudy (anonymous):

That is the complex valued cosine function expanded about the point z = pi / 4. I am not sure of the problem, exactly.

OpenStudy (anonymous):

The radius of convergence is indeed infinite.

OpenStudy (anonymous):

@malevolence19 it's the summed up version, just \[R = \lim_{n \rightarrow infinity} \left| an/an+1 \right|\] where the series is \[\sum_{n=0}^{infinity} an (z-z0)^n\] Z0 being the center of the series. and I meant my center, z0, is supposed to be pi/2, but I get pi/4. I even tried this with a ratio test and still not the right answer

OpenStudy (anonymous):

It is not pi/2. It is pi/4.

OpenStudy (anonymous):

And this theorem is the complex equivalent of the ratio test, FYI.

OpenStudy (anonymous):

Thanks I know, but the correct answer for z0 is pi/2. I don't even know how they get it like that

OpenStudy (anonymous):

No, I am quite sure it is not. Your source may have a typo.

OpenStudy (anonymous):

The textbook? I don't know man

OpenStudy (anonymous):

Here is another one, and now the radius is making no sense. \[\sum_{n=0}^{infinity} (n(n-1))/(2^n) * (z+i)^(2n)\]

OpenStudy (anonymous):

If you are talking about the series \[ \sum_{n=0}^\infty (-1)^n\frac{(z - \frac{\pi}{4})^{2n}}{(2n)!} \] Then the radius of convergence is infinity, the center of the series is pi / 4, and it converges to \[\cos(z-\frac{\pi}{4}) \] I am quite sure of this.

OpenStudy (anonymous):

The radius of convergence I get is \[\sqrt{2}\] and the z0 is -i, however the answer is \[\sqrt{3}\] and i

OpenStudy (anonymous):

And I'm sure of my answer, but I just don't understand how the book's answer is so different.

OpenStudy (anonymous):

Are you sure you're looking at the right answers ... ?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

Well, I'm afraid I can't really give any further input. As far as I can tell, the answers that you gave are correct, and the ones you say are from the book are not. You may of course seek other opinions but I am quite confident that I'm not insane.

OpenStudy (anonymous):

What book are you using, by the way?

OpenStudy (anonymous):

Advanced engineering mathematics by erwin kryszig, international edition

OpenStudy (anonymous):

I found your book, so if you're still around, where exactly are these problems?

OpenStudy (anonymous):

Problem set 15.2, odd questions starting from 7 since they have their answers in the back

OpenStudy (anonymous):

I don't see any of those series, so maybe I have a different edition than you do... What are the page numbers / problem numbers of the two you listed above?

OpenStudy (anonymous):

They are on page 684.. Chapter 15, Section 2

OpenStudy (anonymous):

Mine is tenth edition

OpenStudy (anonymous):

I just looked up the book, and you wrote the series down wrong........

OpenStudy (anonymous):

Read the problems again. I'm glad this got settled and I'm not crazy, but please in the future make sure that you read the problem correctly... :)

OpenStudy (anonymous):

How? 7. \[\sum_{n=0}^{infinity} ((-1)^n)/(2n!)) (z-1/4 \pi)^(2n)\] I don't think I wrote it wrong..

OpenStudy (anonymous):

If I did please tell me because this is driving me insane

OpenStudy (anonymous):

Oh bloody flutterers. The answers are to the questions you've put up. They didn't update the answers to the changed questions

OpenStudy (anonymous):

Can I have the link to that book?

OpenStudy (anonymous):

You may indeed http://megaupload.com/?d=1GND4DET

OpenStudy (anonymous):

Thank you, now I can actually do the real questions.